# CREME: Robustness Enhancement of Code LLMs via Layer-Aware Model Editing

### v2

@@ 7.2. Knowledge Editing
Knowledge editing has emerged as a promising direction for updating factual knowledge in LLMs without full retraining (Zhong et al., 2023; Belrose et al., 2023; Wei et al., 2023; Gupta et al., 2024; Hase et al., 2023a; Hua et al., 2024; Lo et al., 2024), which can be categorized into three main paradigms (Wang et al., 2024c). External memorization-based methods (Madaan et al., 2022; Zheng et al., 2023; Murty et al., 2022; Mitchell et al., 2022; Li et al., 2022; Dong et al., 2022; Huang et al., 2023; Hartvigsen et al., 2023; Li et al., 2024a, b) leverage an external memory to store new knowledge without modifying the pre-trained weights, thereby fully preserving the original knowledge encoded in the LLM. Global optimization-based methods (Chen et al., 2020; Sinitsin et al., 2020; Lee et al., 2022; Zhu et al., 2020; Ni et al., 2023; Gu et al., 2024; De Cao et al., 2021; Hase et al., 2023b) aim to integrate new knowledge into pre-trained LLMs in a generalizable manner through optimization guided by the new information. These methods employ tailored strategies to constrain the impact on existing knowledge, distinguishing them from naive fine-tuning. The third category, local modification-based methods (Dai et al., 2021; Wu et al., 2023; Li et al., 2024c), is most relevant to our work. These methods aim to identify the parameters associated with specific knowledge in LLMs and selectively update them to incorporate new information related to the edit.
Meng et al.(Meng et al., 2022a) propose ROME, a pioneering local editing method that identifies and updates key neuron activations within specific transformer layers responsible for storing factual associations. By analytically tracing causal dependencies, ROME directly modifies the internal representations of a model to reflect new factual information with high locality and minimal side effects. Building on this, Meng et al.(Meng et al., 2022b) introduce MEMIT, a scalable framework that enables batch editing of multiple facts by optimizing a shared intervention across selected layers. Wang et al. (Wang et al., 2024b) extend this line of work to the domain of safety, applying knowledge editing to mitigate toxic behaviors in LLMs by targeting offensive content and aligning representations with non-toxic alternatives. These methods underscore the potential of local interventions to precisely and efficiently influence LLM behavior. However, prior approaches have primarily focused on factual updates, whereas our work explores the novel application of local knowledge editing to improve robustness against prompt perturbations.

### v3

@@ 7.2. Knowledge Editing
Knowledge editing has emerged as a promising direction for updating factual knowledge in LLMs without full retraining (zhong2023mquake; belrose2023leace; wei2023assessing; gupta2024model; hase2023does; hua2024propagation; lo2024large), which can be categorized into three main paradigms (wang2024knowledge). External memorization-based methods (madaan2022memory; zheng2023can; murty2022fixing; mitchell2022memory; li2022large; dong2022calibrating; huang2023transformer; hartvigsen2023aging; li2024consecutive; li2024swea) leverage an external memory to store new knowledge without modifying the pre-trained weights, thereby fully preserving the original knowledge encoded in the LLM. Global optimization-based methods (chen2020recall; sinitsin2020editable; lee2022plug; zhu2020modifying; ni2023forgetting; gu2024model; de2021editing; hase2023methods) aim to integrate new knowledge into pre-trained LLMs in a generalizable manner through optimization guided by the new information. These methods employ tailored strategies to constrain the impact on existing knowledge, distinguishing them from naive fine-tuning. The third category, local modification-based methods (dai2021knowledge; wu2023depn; li2024pmet), is most relevant to our work. These methods aim to identify the parameters associated with specific knowledge in LLMs and selectively update them to incorporate new information related to the edit.
Meng et al.(meng2022locating) propose ROME, a pioneering local editing method that identifies and updates key neuron activations within specific transformer layers responsible for storing factual associations. By analytically tracing causal dependencies, ROME directly modifies the internal representations of a model to reflect new factual information with high locality and minimal side effects. Building on this, Meng et al.(meng2022mass) introduce MEMIT, a scalable framework that enables batch editing of multiple facts by optimizing a shared intervention across selected layers. Wang et al. (wang2024detoxifying) extend this line of work to the domain of safety, applying knowledge editing to mitigate toxic behaviors in LLMs by targeting offensive content and aligning representations with non-toxic alternatives. These methods underscore the potential of local interventions to precisely and efficiently influence LLM behavior. However, prior approaches have primarily focused on factual updates, whereas our work explores the novel application of local knowledge editing to improve robustness against prompt perturbations.